nini:

Sam buys one bag of flour for $2.50, three packages of shredded cheese for $3.00 each, and x packages of mushrooms for $3.50 each. If Sam paid $18.50 in all, how many packages of mushrooms did Sam buy? 4 3 1 2

nini:

@darkknight @Imagine @mxddi3

mxddi3:

well, actually it should be 2.50+3(3.00)+3.50x=18.50

nini:

but i dont get it still kindaaaaa....

darkknight:

Yes maddie, that is the correct formula, just saying that so that you don't get confused nini

nini:

yup still dont get it

mxddi3:

ok so we have 2.50+3(3.00)+3.50x=18.50 do you know where i got those numbers from?

Imagine:

\[ 11.50+3.50x=18.50\]

nini:

no mxddiex

nini:

oh so is it 1

mxddi3:

ok so one bag of flour is 2.50 and he got 3 packages of shredded cheese for $3 each and then the mushrooms are 3.50 but we dont know how many they got, so we have 3.50x. and the total he spends is 18.50 ok so then we combine like terms, as Imagine did above. do you get this so far?

Imagine:

so he bought one flour which equals to 2.50. he bought 3 peices of shredded cheese, which equals 9.00. Then it is 3.50 per each mushroom, we are trying to figure out how much mushrooms he got.

nini:

so it s 3

Imagine:

so the total adds to 18.50

Imagine:

incorrect.

mxddi3:

if we have 11.50+3.50x=18.50 the next step is to subtract 11.50 from both sides. so on the left we have 3.50x and on the right we have 18.50-11.50=7 so we now have 3.50x=7 can you do this for me please nini?